Home‎ > ‎Crew 160‎ > ‎

0107-DailySummaryReport

MDRS Daily Summary Report for sol 12 1/7/2016

And More Star Wars

Executive Commander Reporting

Mission Status: Today the crew had an eventful EVA and enjoyed crew resting period.

Sol Activity Summary: Some of the crew woke up early and some didn’t, but the crew still enjoyed a nice loving breakfast that Commander whipped up. After breakfast the crew decided to jump steadfast into today’s EVA to Piccard’s Butte. The EVA was going well till one of crew members said there was a storm brewing. Turns out the sky lied to us as soon as everyone came back end. All is well for the rest of the day since we decided to continue with our Star Wars marathon and plan to finish by the end of the rotation.

Look Ahead Plan: Crew is still working hard on finishing up Star Wars. Research is continuing as well.

Anomalies in work: There are no anomalies to report.

Weather: The Martian sky was cloudy with somewhat chance of rain?

Crew Physical Status: Each crew member is doing fine!

EVA: 12S-052183E-425130N

Reports to be filed:

Commanders Report

Journalist Report

GreenHab Report

Engineering Report

EVA #11 Summary Report

Support Requested: N/A